Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probability and Combination Problems

Problem Solving — algebra and arithmetic
Expert replies

by vsr2014 » Tue Oct 21, 2014 9:59 am
beatthegmat wrote:Found this great set of probability and combination problems. Remember--these types of questions appear infrequently on the GMAT, so don't over-emphasize them too much in your studies!
#48 The solution provided doesn't seem to answer the right question. The question in the problem is "In order to ensure that the sum of all cards he drew was even, how many cards did Jerome have to draw?"
The solution says: The 12th draw ensures an even sum.
Let's assume Jerome drew 12 cards and there are 9 odd cards and 2 even cards, so the sum is ODD
If he drew 13 cards, 9 odds and 4 evens, the sum is still ODD.
SO he has to draw all 20 cards to ensure that the sum is EVEN.
I assume that he is not calculating the sum upon drawing every next card, because the condition doesn't say so..
Is my logic wrong here?
Join the discussion

by [email protected] » Tue Oct 21, 2014 10:53 am
Hi tbasebal24 and vsr2014,

The original post is over 2 years old and the response appears to be to a different question, but for anyone else who might be interested… here's how you can solve this problem.

In probability questions, you can either calculate what you "want" or what you "don't want"; certain probability questions are easier to deal with if you focus on what you "don't want" (since the two possibilities will always add up to 1).

Here, we have 3 red marbles and 7 blue marbles. We're told to select 2 marbles at random and we're asked for the probability that AT LEAST ONE will be blue. Instead of calculating all the different ways to get 1 or 2 blue marbles, let's calculate what we DON'T want: 0 blue marbles.

On the first pull, we'd have a 3/10 chance of NOT getting a blue marble. After pulling out that first marble….
On the second pull, we'd have a 2/9 chance of NOT getting a blue marble.

(3/10)(2/9) = 6/90 = 1/15 --> this is the probability of NOT getting what we want.

1 - 1/15 = 14/15 --> this is the probability of getting what we WANT.

Final Answer: D

GMAT assassins aren't born, they're made,
Rich
Contact Rich at [email protected]
Image
Join the discussion